An employee of a big business company earns a gross annual salary of



₱450,000. What is the weekly, and monthly pay for this employee?


1
Expert's answer
2022-04-18T16:45:14-0400

"Weekly:\\450000\cdot \frac{7}{365}=8630.14\\Monthly:\\450000\cdot \frac{1}{12}=37500"